. A piano string of length 76cm and mass per unit length of 6.91 x 10-3
kg/m is under a
tension of value 40N. What is the frequency of a note sounded by this string? (2 marks
1
Expert's answer
2020-07-01T15:42:11-0400

The frequency of a string in terms of the given quantities can be calculated as


"f=\sqrt{\frac{T}{4L^2\mu}}=\sqrt{\frac{40}{4\cdot0.76^2\cdot6.91\cdot10^{-3}}}=51\text{ Hz}."